Unpolarized light falls on two polarizing sheets placed one on top of the other. What must be the angle





between their transmission axis, if intensity of transmitted light by the analyzer is one third of that of





incident light?






1
Expert's answer
2022-01-19T14:52:01-0500

"I_2=I_1cos^2\theta"

"\frac{1}{3}I_1=I_1cos^2\theta\implies \theta=cos^{-1}\frac{1}{√{3}}"

"\therefore \theta=55°"